what are the last two digits of 2 raised to the power of 64?please give me the solution too.thanks.

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

One can observe

Last 2 digits of

24^even is 76

24^odd is 24

So

2^64=2^60*2^4

=(1024)^6*2^4

So last 2 digits are 76*16=1216

So 16 is the answer

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..